3y=x-1 and 6y=x find a common solution for these two equations
Answer by gonzo(654)

3y = x - 1
6y = x
can substitute 6y for x in the first equation.
so 3y = 6y - 1
-3y = -1
y = -1/-3 = 1/3.
using that value solve for x in either equation.
in the second equation, 6 * (1/3) = 2 = x, so x = 2.
in the first equation, 3 * (1/3) = 1 = x - 1, so x = 2 again (as it should).
answer is x = 2 and y = 1/3.